Evaluate the moment of inertia of a tin rod with axis of rotation at the center ,if its lenght is 100cm and its mass is 2.3kg



What is the difference between Rotational K.E And linear K.E



write down tgere formular and compare them

1
Expert's answer
2022-01-19T10:12:07-0500

Solution

Mass of tin rod M =2.3Kg

It's length L=1m

the moment of inertia of a tin rod with axis of rotation at the center is given by

"I=mL^2\\=2.3\times(1) ^2
\\=2.3kg-m^2"

Linear kinetic energy is given by

"KE=\frac{mv^2}{2}"

Rotational kinetic energy

"I=\frac{I\omega^2}{2}"


These are interconnected.